OpenStudy (anonymous):

A. 8 B. 36 C. 4.5 D. 9

OpenStudy (anonymous):

what is the solution to the problem 6/x= 9/12

OpenStudy (anonymous):

@Graeze_nanquil

OpenStudy (anonymous):

Oh, okay. Thank you!

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

i think it is d

OpenStudy (campbell_st):

nope its not D start by cross multiplying... 6*12 = 9x

OpenStudy (anonymous):

6 x 12 = 72

OpenStudy (campbell_st):

yes, so you need to solve 72 = 9x

OpenStudy (anonymous):

9 x 8 = 72